# Learning About Social Context from Smartphone Data: Generalization Across Countries and Daily Life Moments

### 29 Jan 2024 v4

@@ Abstract.
mobile sensing, social context, smartphone sensors, machine learning, context-awareness
journalyear: 2024copyright: acmlicensedconference: Proceedings of the CHI Conference on Human Factors in Computing Systems; May 11–16, 2024; Honolulu, HI, USAbooktitle: Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24), May 11–16, 2024, Honolulu, HI, USAdoi: 10.1145/3613904.3642444isbn: 979-8-4007-0330-0/24/05
@@ 1. Introduction
Human beings inherently need social interactions, community, and interpersonal relationships (Baumeister and Leary, 1995). A fulfilling social life is particularly important for the mental well-being of young adults, as studies have shown (Fuligni and Eccles, 1993; Weeks and Asher, 2012; Blakemore, 2012). Regular interaction with friends, family, or colleagues during adolescence positively impacts long-term mental and physical health. Conversely, social isolation, such as living alone or prolonged solitude, increases the likelihood of depression and other mental and physical health issues (Holt-Lunstad et al., 2015; Matthews et al., 2016; Vanhalst et al., 2015; Baumeister and Leary, 1995; Blakemore, 2012). Therefore, understanding an individual’s social context over time is beneficial in mobile health applications, enabling timely interventions and feedback (Meegahapola et al., 2020a, 2021a). Beyond that, "a system is context-aware if it uses context to provide relevant information and/or services to the user, where relevancy depends on the user’s task" (Dey, 2001), and social context is a key attribute in making mobile applications context-aware, enhancing their relevance and effectiveness (Raento et al., 2005; Liang and Cao, 2015; Chen and Kotz, 2000; Dey et al., 2001; Porfirio et al., 2020; Mayer et al., 2015). Hence, inferring social context could also benefit context-aware computing, enabling devices and applications, like smartphones, to adapt functionalities (i.e., displaying notifications, silencing the phone, recommendations, etc.) based on the user’s social setting. Hence, this inference not only improves user engagement through personalized interactions, but also aids digital health applications to provide timely support and enhance the overall well-being of users.
@@
  • Contribution 4: In the specific cases of mood inference and complex daily activity recognition, even though prior work found that models might generalize reasonably to geographically closer countries in Europe, we did not find such associations in social context inference performance. In fact, in all cases, models did not generalize well to unseen countries regardless of the geographical proximity. This raises concerns regarding where trained social context inference models could be deployed.
@@
Term Description
Social Context
If a person is Alone or Not (i.e., with friends, family, or others). We borrow this definition from prior studies in human-computer interaction and ubiquitous computing domain (Burns et al., 2011; Meegahapola et al., 2021a; Kammoun et al., 2023; Meegahapola et al., 2020a).
Country-Specific Approach
Machine learning models are trained and tested only on data from one specific country.
Country-Agnostic Approach
Machine learning models are trained on one country or a set of countries and tested on another country not seen in training. Such models are usually trained with the assumption that the model could be deployed to other countries, hence the term agnostic.
Multi-Country Approach
Machine learning models are trained on a set of countries and tested again on the same set of countries. This is the generic way of training models when data from multiple countries are available.
Population-Level Model (PLM)
Training and Testing splits have a disjoint set of users. Represents a case where a machine learning model trained with a population is deployed to a mobile app that a new user uses. Hence, end-user data have not been used in model training, leading to non-personalized and generic one-size-fits-all models.
Hybrid Model (HM)
Training and testing splits do not have a disjoint set of users. Represent a case when a mobile app user uses a machine learning model for some time, and data from the user is used in re-training (or fine-tuning in the case of neural networks) models. Hence, this approach leads to partially personalized models.
@@ 3. Dataset
Table 2. Data summary statistics of the used dataset
country # of participants # of participants with sufficient data # of reports alone percentage
UK
72
53
26,687 69.05%
Denmark
25
17
10,058 49.63%
Italy
238
221
151,335 64.78%
Paraguay
29
24
9,745 54.52%
Mongolia
217
138
94,249 24.52%
All
581
453
292,074 51.31%
We obtained a novel smartphone sensing dataset that was collected in a study conducted simultaneously in five countries, from our previous work (Assi et al., 2023; Giunchiglia et al., 2022). To the best of our knowledge, this is the only dataset that allows us to perform the analysis around social context during daily life moments, across multiple countries, with smartphone data. The data were collected for four weeks from college students of the following five universities: Aalborg University (Denmark), London School of Economics and Political Science (the United Kingdom), the National University of Mongolia (Mongolia), the Universidad Católica "Nuestra Señora de la Asunción" (Paraguay), and the University of Trento (Italy). The study participants contributed three different kinds of data: (i) closed-ended questionnaires, (ii) hourly self-reports throughout the day, and (iii) sensor data.
@@
The final analysis of this study consists of inferring the social context of a participant. Given the large size of the original dataset, which amounts to approximately 30GB of data in different data formats, the sensor data were processed and aggregated to sensible features at the self-report level. To do this, we used an approach similar to prior mobile sensing studies where sensor data are aggregated around the time of an in-situ self-report (Servia-Rodríguez et al., 2017; Bae et al., 2017; Biel et al., 2018; Meegahapola et al., 2021a, 2020a). More specifically, all sensor data corresponding to ten minutes around the social context self-report was aggregated to features as summarized in Table 6. Thus, the aggregation time frame corresponds to a ten-minute window around the answering time of a self-report, which captures the individual’s current activity, semantic location, mood, and, finally, the social context. It is worth noting that various time windows ranging from 1 to 20 minutes were explored. Extending beyond a 20-minute window was not feasible due to the risk of data leakage from overlapping sensor data segments. Among the tested durations, the ten-minute window emerged as reasonable, offering satisfactory performance. Additionally, this specific time frame aligns with previous studies that used the same dataset (Meegahapola et al., 2023; Assi et al., 2023), thereby facilitating meaningful comparisons. Moreover, in total, 117 distinct sensor features were created in line with the work of the original authors (Assi et al., 2023), including 48 features that correspond to app usage time period based on app categories in Google Playstore (e.g., action, dating, music, puzzle, social, etc.) similar to (Santani et al., 2018; LiKamWa et al., 2013). Additionally, time and day features were also included in the analysis, where the day of the week corresponds to a categorical variable, weekday or weekend, while the time was noted with the hour of the corresponding time dairy event (a numeric value between 0 and 23). Furthermore, missing data markers were introduced, which indicate when data is missing for certain sensor groups (Little and Rubin, 2019). In summary, the smartphone sensor features of the final dataset used for inference are detailed in Appendix A.
@@
Figure 1. Ratio of missing data (Nan values) in percentage aggregated to sensor group
Figure 1. Ratio of missing data (Nan values) in percentage aggregated to sensor group
@@
Figure 2. Boxplot of the number of self-reports per participant, grouped by country. The boxes extend till the upper and lower quartiles, the median is indicated by an orange line, and whiskers extend till 1.5x IQR
@@ 4.1. RQ1: Situational and Contextual Aspects Around Social Contexts
The purpose of this analysis was to investigate various situational and behavioral cues related to social context (alone or not) that can be extracted from self-reports and sensor data. This examination provides us with a comprehensive overview of the dataset and positions our study in relation to prior research. Initially, the original dataset featured detailed social context categories, including labels such as alone, with classmates, with colleagues, with a partner, with relatives, with roommates, and with others. Utilizing these labels, we examined how patterns of fine-grained social contexts differed across countries (Figure 3). Subsequently, by aggregating data from all countries, we explored how social context varied across these fine-grained categories throughout the day (Figure 4). Next, as previous studies have linked social context to mood (discussed in Section 2), we conducted an analysis focusing on binary social context (i.e., alone or not) across different levels of valence, assessed using a five-point Likert scale ranging from very positive to very negative. Lastly, we delved into the variations in social context across different locations and activities (Figure 5).
@@ 4.2. RQ2, RQ3 & RQ4: Social Context Inference
  • Hybrid models (partially personalized): First, data in training and testing splits are not from the same users, $\approx$ 60:40 split. Then, add 50% of data from testing users into the training set to achieve partial personalization for users in the testing set, leading to $\approx$ 80:20 training and testing split. Moreover, temporally, only the first half of data produced by testing users are removed from the testing set and included in the training set to make sure there is no temporal leakage (Kapoor and Narayanan, 2022). That is, no data points from the future of the testing users are used in model training.
@@
Furthermore, behavioral sensor data often exhibit sparsity and tend to contain high volumes of missing data (Xu et al., 2021). Within this dataset, missing data can be attributed to various distinct scenarios: (i) sensor malfunction, (ii) sensor damage or absence, and (iii) intentional sensor deactivation by the user (e.g., disabling Bluetooth/WiFi, activating flight mode, etc.). However, distinguishing between these cases based solely on the available sensor data is challenging, and as in previous studies (Pires et al., 2016; Stisen et al., 2015), all these scenarios are treated equivalently. The extent of missing data also varies depending on the sensor (Figure 2) and the country where the experiment was conducted (e.g., 60% of Mongolian sensor data is missing). To address the issue of missing data, the first step involved discarding sensor groups with over 90% missing data, aligning with prior work (Santani et al., 2018). This exclusion specifically applied to the ’cellular_gsm’ features (refer to Figure 2). Subsequently, the remaining missing observations were estimated using the k-nearest-neighbor (kNN) algorithm (Zhou et al., 2018; Xu et al., 2021), with an experimentally determined optimal value of two for k. Following the procedure outlined in prior research (Assi et al., 2023; Zhou et al., 2018; Xu et al., 2021), estimation was performed on the training set, and the learned model was then employed to impute missing data in the testing sets. To retain information that might be encoded in the missing entries of a given sensor group, binary markers were used to denote that the data had been interpolated (Little and Rubin, 2019).
@@ 5.1. RQ1: Situational and Contextual Aspects Around Social Contexts
Figure 5 depicts the co-occurrence of specific moods (positive and negative valence), activities, and semantic locations with the social context (alone or not). Regarding concurrent activities and whether someone is alone or not, activities such as studying, listening to lectures, sleeping, and personal care were predominantly carried out alone. In contrast, activities involving social interaction, such as family care, eating, and nurturing one’s social life, were typically done with others. Furthermore, regarding semantic location, participants reported having been alone at home on 125,270 instances, which could be influenced by the COVID-19 pandemic and the prevalence of remote work and study environments (Fried et al., 2022). However, given the remote work culture that is prevalent even after the pandemic, we could expect these patterns to hold to an extent (Lund et al., 2021; Broom, 2021; Mass, 2021). In terms of gender-specific differences, there were no notable disparities between men and women. Men reported being alone in 54.6% of all self-reports, while women reported to be slightly less alone, accounting for 49.1% of all self-reports. Additionally, women reported to have been with their families, averaging 31.6%, which is approximately 5% more than men. All other distinctions between women and men were minor in the broader context. These findings align with existing research indicating that young men tend to experience more loneliness (Barreto et al., 2021) and tend to be alone more in certain contexts (Fry and Parker, 2021). Prior work also discussed similar findings that women tend to seek more social ties compared to men (Tamminen et al., 2019). However, it is important to note that being alone could be associated with a wide range of factors, including culture, geography, age, etc., in addition to gender (Barreto et al., 2021).
@@
feature coefficient std. error z-score p-value feature coefficient std. error z-score p-value
Denmark app_news_and_magazines -0.105 0.041 -2.537 0.011 Paraguay app_finance -0.043 0.012 -3.468 0.000
app_casual -0.041 0.016 -2.492 0.012 notifications_w/o_duplicates -0.018 0.005 -3.248 0.001
app_simulation -0.180 0.074 -2.416 0.015 app_productivity -0.000 $0$3.195 0.001
app_action -0.078 0.032 -0.241 0.015 location_altitude -0.000 $0$3.175 0.001
app_books_and_reference -2.651 1.099 -2.412 0.015 activity_in_vehicle -0.004 0.001 -2.807 0.004
app_sports -0.582 0.241 -2.412 0.015 activity_running -0.025 0.010 -2.588 0.009
app_adventure -0.109 0.045 -2.409 0.015 notifications_removed -0.003 0.001 -2.577 0.009
proximity_std -0.027 0.011 -2.375 0.017 wifi_number_of_devices -0.007 0.003 -2.466 0.013
app_business -0.017 0.007 -2.364 0.018 activity_on_bicycle -0.016 0.006 -2.405 0.016
app_maps_and_navigation -0.352 0.149 -2.354 0.018 wifi_minimum_rssi -0.006 0.002 -2.172 0.029
UK user_presense_time -0.000 0.000 1.987 0.046 Mongolia app_video_players_editors 0.000 0.000 2.613 0.008
proximity_minimum -0.008 0.004 -1.975 0.048 app_shopping -0.056 0.021 -2.609 0.009
proximity_mean -0.018 0.009 -1.909 0.056 app_travel_and_local -0.056 0.022 -2.500 0.012
screen_number_of_episodes 0.011 0.006 -1.844 0.065 app_trivia -0.674 0.284 -2.373 0.017
app_photography 0.006 0.003 1.802 0.701 app_books_and_reference -0.016 0.007 -2.312 0.020
steps_counter 0.000 0.000 1.775 0.075 app_educational -0.109 0.047 -2.300 0.021
proximity_maximum -0.015 0.009 -1.608 0.107 app_weather -3.481 1.516 -2.295 0.021
app_entertainment 0.020 0.012 1.588 0.112 app_comics -0.379 0.165 -2.291 0.021
activity_on_bicycle 0.007 0.004 1.559 0.118 app_art_and_design -0.705 0.315 -2.238 0.025
app_business 0.003 0.002 1.440 0.149 app_news_and_magazines -0.134 0.060 -2.227 0.025
Italy activity_in_vehicle 0.004 0.000 5.970 0.000 All app_news_and_magazines -0.105 0.041 -2.537 0.011
activity_onfoot 0.003 0.000 4.092 0.000 app_casual -0.041 0.016 -2.492 0.012
activity_walking 0.003 0.000 4.010 0.000 app_simulation -0.180 0.074 -2.416 0.015
steps_counter 0.000 0.000 3.243 0.001 app_action -0.078 0.032 -2.414 0.015
activity_tilting 0.000 0.000 3.225 0.001 app_books_and_reference -2.651 1.099 -2.412 0.015
location_altitude 0.000 0.000 3.164 0.001 app_sports -0.582 0.241 -2.412 0.015
proximity_mean -0.013 0.004 -3.067 0.002 app_adeventure -0.109 0.045 -2.409 0.015
notifications_w/o_duplicates -0.006 0.002 -2.409 0.015 proximity_std -0.027 0.011 -2.375 0.017
wifi_std_rssi 0.004 0.001 2.394 0.016 app_business -0.017 0.007 -2.364 0.018
screen_number_of_episodes 0.003 0.001 2.111 0.034 app_maps_and_navigation -0.352 0.149 -2.354 0.018
@@
The statistical analysis presented in Table 3 highlights the top ten sensing features effective in discerning whether a person is alone or not, based on the p-value after Bonferroni correction. This analysis accounts for both fixed (features) and random effects (study participants), offering a nuanced understanding of the data. In Denmark, app usage features such as news and magazines, simulation, and books and references show moderate to high negative coefficients, indicating that using these apps is inversely related to being with others. Conversely, in Paraguay, app usage features like finance and productivity, along with the number of notifications received, have smaller negative coefficients. In the UK, features related to user presence and proximity are statistically significant but show almost no relationship with social context. Similarly, in Mongolia, features like trivia, comics, weather, and arts and design have notable negative relationships with being alone. Italy displays a unique pattern where physical activities such as being in a vehicle, on foot, and walking, though statistically significant, have very small coefficients, showing no clear association. This pattern, at least in terms of feature set, is akin to the UK and Paraguay but contrasts with Denmark and Mongolia, where app usage features have the top statistical significance values. It is also worth noting that, in the UK, except for two features (user_presense_time and proximity_minimum), other features had $p\geq 0.05$, suggesting no statistical significance. The ’All’ category, aggregating data from all countries, indicates that app features like news and magazines, books and references, maps, sports, and simulation are generally negatively associated with being alone. These findings highlight country-specific differences in the relationship between sensor data features and social context, emphasizing the need to consider user behavior diversity in different countries for understanding and inferring everyday life behaviors with multimodal sensor data (Khwaja et al., 2019; Assi et al., 2023).

### 01 Mar 2024 v5

@@ Abstract.
mobile sensing, smartphone sensing, multimodal, social context, machine learning, generalization, personalization, context-awareness, digital health
journalyear: 2024copyright: acmlicensedconference: Proceedings of the CHI Conference on Human Factors in Computing Systems; May 11–16, 2024; Honolulu, HI, USAbooktitle: Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24), May 11–16, 2024, Honolulu, HI, USAdoi: 10.1145/3613904.3642444isbn: 979-8-4007-0330-0/24/05ccs: Human-centered computing Empirical studies in ubiquitous and mobile computingccs: Human-centered computing Smartphonesccs: Human-centered computing Empirical studies in HCIccs: Social and professional topics Geographic characteristicsccs: Social and professional topics Cultural characteristics
@@ 1. Introduction
Human beings inherently need social interactions, community, and interpersonal relationships (Baumeister and Leary, 1995). As studies have shown, a fulfilling social life is particularly important for the mental well-being of young adults (Fuligni and Eccles, 1993; Weeks and Asher, 2012; Blakemore, 2012). Regular interaction with friends, family, or colleagues during adolescence positively impacts long-term mental and physical health. Conversely, social isolation, such as living alone or prolonged solitude, increases the likelihood of depression and other mental and physical health issues (Holt-Lunstad et al., 2015; Matthews et al., 2016; Vanhalst et al., 2015; Baumeister and Leary, 1995; Blakemore, 2012). Therefore, understanding an individual’s social context over time is beneficial in mobile health applications, enabling timely interventions and feedback (Meegahapola et al., 2020a, 2021a). Beyond that, "a system is context-aware if it uses context to provide relevant information and/or services to the user, where relevancy depends on the user’s task" (Dey, 2001), and social context is a key attribute in making mobile applications context-aware, enhancing their relevance and effectiveness (Raento et al., 2005; Liang and Cao, 2015; Chen and Kotz, 2000; Dey et al., 2001; Porfirio et al., 2020; Mayer et al., 2015). Hence, inferring social context could also benefit context-aware computing, enabling devices and applications, like smartphones, to adapt functionalities (i.e., displaying notifications, silencing the phone, recommendations, etc.) based on the user’s social setting. Hence, this inference not only improves user engagement through personalized interactions, but also aids digital health applications to provide timely support and enhance the overall well-being of users.
@@
  • Contribution 4: In the specific cases of mood inference and complex daily activity recognition, even though prior work found that models might generalize reasonably to geographically closer countries in Europe, we did not find such associations in social context inference performance. In fact, in all cases, models did not generalize well to unseen countries regardless of the geographical proximity. This raises concerns regarding where a trained social context inference models could be deployed.
@@
[…]
This table provides terminology and definitions used in this paper, including terms like social context, country-specific approach, country-agnostic approach, multi-country approach, population-level model, and hybrid model. Term Description Social Context If a person is Alone or Not (i.e., with friends, family, or others). We borrow this definition from prior studies in human-computer interaction and ubiquitous computing domain (Burns et al., 2011; Meegahapola et al., 2021a; Kammoun et al., 2023; Meegahapola et al., 2020a). Country-Specific Approach Machine learning models are trained and tested only on data from one specific country. Country-Agnostic Approach Machine learning models are trained on one country or a set of countries and tested on another country not seen in training. Such models are usually trained with the assumption that the model could be deployed to other countries, hence the term agnostic. Multi-Country Approach Machine learning models are trained on a set of countries and tested again on the same set of countries. This is the generic way of training models when data from multiple countries are available. Population-Level Model (PLM) Training and Testing splits have a disjoint set of users. Represents a case where a machine learning model trained with a population is deployed to a mobile app that a new user uses. Hence, end-user data have not been used in model training, leading to non-personalized and generic one-size-fits-all models. Hybrid Model (HM) Training and testing splits do not have a disjoint set of users. Represent a case when a mobile app user uses a machine learning model for some time, and data from the user is used in re-training (or fine-tuning in the case of neural networks) models. Hence, this approach leads to partially personalized models.
@@ 3. Dataset
Table 2. Dataset Summary.
[…]
Summary statistics of the dataset used in this study are presented, covering five countries: Denmark, Italy, Mongolia, Paraguay, and the UK. The paper uses data from 581 participants who provided over 292,000 self-reports on their social context, with a roughly equal distribution between ’alone’ (around 51%) and ’not alone’ cases. Country # of Participants # of Participants With Sufficient Data # of Reports Alone Percentage UK 72 53 26,687 69.05% Denmark 25 17 10,058 49.63% Italy 238 221 151,335 64.78% Paraguay 29 24 9,745 54.52% Mongolia 217 138 94,249 24.52% All 581 453 292,074 51.31%
For this study, we used a novel smartphone sensing dataset from our previous work, that was collected in a study conducted simultaneously in five countries (Assi et al., 2023; Giunchiglia et al., 2022). To the best of our knowledge, this is the only dataset that allows us to perform the analysis around social context during daily life moments, across multiple countries, with smartphone data. The data were collected for four weeks from college students of the following five universities: Aalborg University (Denmark), London School of Economics and Political Science (the United Kingdom), the National University of Mongolia (Mongolia), the Universidad Católica "Nuestra Señora de la Asunción" (Paraguay), and the University of Trento (Italy). The study participants contributed three different kinds of data: (i) closed-ended questionnaires, (ii) hourly self-reports throughout the day, and (iii) sensor data.
@@
The final analysis of this study consists of inferring the social context of a participant. Given the large size of the original dataset, which amounts to approximately 30GB of data in different data formats, the sensor data were processed and aggregated to sensible features at the self-report level. To do this, we used an approach similar to prior mobile sensing studies where sensor data are aggregated around the time of an in-situ self-report (Servia-Rodríguez et al., 2017; Bae et al., 2017; Biel et al., 2018; Meegahapola et al., 2021a, 2020a). More specifically, all sensor data corresponding to ten minutes around the social context self-report was aggregated to features as summarized in Table 6. Thus, the aggregation time frame corresponds to a ten-minute window around the answering time of a self-report, which captures the individual’s current activity, semantic location, mood, and, finally, the social context. It is worth noting that various time windows ranging from 1 to 20 minutes were explored. Extending beyond a 20-minute window was not feasible due to the risk of data leakage from overlapping sensor data segments. Among the tested durations, the ten-minute window emerged as reasonable, offering satisfactory performance. Additionally, this specific time frame aligns with previous studies that used the same dataset (Meegahapola et al., 2023; Assi et al., 2023), thereby facilitating meaningful comparisons. Moreover, in total, 117 distinct sensor features were created in line with the work of the original authors (Assi et al., 2023), including 48 features that correspond to app usage time period based on app categories in Google Playstore (e.g., action, dating, music, puzzle, social, etc.) similar to (Santani et al., 2018; LiKamWa et al., 2013). Additionally, time and day features were also included in the analysis, where the day of the week corresponds to a categorical variable, weekday or weekend, while the time was noted with the hour of the corresponding time dairy event (a numeric value between 0 and 23). Furthermore, missing data markers were introduced, which indicate when data is missing for certain sensor groups (Little and Rubin, 2019). In summary, the smartphone sensor features of the final dataset used for inference are detailed in Table 6 in the Appendix.
@@
Figure 1. Ratio of missing data (Nan values) in percentage aggregated to sensor group.
@@
Figure 2. Boxplot of the number of self-reports per participant, grouped by country. The boxes extend till the upper and lower quartiles, the median is indicated by an orange line, and whiskers extend till 1.5x IQR.
This figure illustrates the percentage of missing data (NaN values) aggregated by the sensor group. It shows that cellular signals have the highest missing values, while touch events have the lowest.
A boxplot representing the number of self-reports per user, categorized by country. The boxes indicate the upper and lower quartiles, with an orange line for the median and whiskers extending to 1.5 times the interquartile range. Italy generally has the highest per-person self-reports, while Denmark has the lowest.
Figure 1. Ratio of missing data (Nan values) in percentage aggregated to sensor group.
Figure 2. Boxplot of the number of self-reports per participant, grouped by country. The boxes extend till the upper and lower quartiles, the median is indicated by an orange line, and whiskers extend till 1.5x IQR.
@@ 4.1. RQ1: Situational and Contextual Aspects Around Social Contexts
The purpose of this analysis was to investigate various situational and behavioral cues related to social context (alone or not) that can be extracted from self-reports and sensor data. This examination provides us with a comprehensive overview of the dataset and positions our study in relation to prior research. Initially, the original dataset featured detailed social context categories, including labels such as alone, with classmates, with colleagues, with a partner, with relatives, with roommates, and with others. Utilizing these labels, we examined how patterns of fine-grained social contexts differed across countries (Figure 3). Subsequently, by aggregating data from all countries, we explored how social context varied across these fine-grained categories throughout the day (Figure 4). Next, as previous studies have linked social context to mood (discussed in Section 2), we conducted an analysis focusing on binary social context (i.e., alone or not) across different levels of valence, assessed using a five-point Likert scale ranging from very positive to very negative. Lastly, we delved into the variations in social context across different locations and activities (Figure 5c).
@@ 4.2. RQ2, RQ3 & RQ4: Social Context Inference
  • Hybrid models (partially personalized): First, data in training and testing splits are not from the same users, and we did an $\approx$ 60:40 split. Then, add 50% of data from testing users into the training set to achieve partial personalization for users in the testing set, leading to $\approx$ 80:20 training and testing split. Moreover, temporally, only the first half of data produced by testing users are removed from the testing set and included in the training set to make sure there is no temporal leakage (Kapoor and Narayanan, 2022). That is, no data points from the future of the testing users are used in model training.
@@
Furthermore, behavioral sensor data often exhibit sparsity and tend to contain high volumes of missing data (Xu et al., 2021). Within this dataset, missing data can be attributed to various distinct scenarios: (i) sensor malfunction, (ii) sensor damage or absence, and (iii) intentional sensor deactivation by the user (e.g., disabling Bluetooth/WiFi, activating flight mode, etc.). However, distinguishing between these cases based solely on the available sensor data is challenging, and as in previous studies (Pires et al., 2016; Stisen et al., 2015), all these scenarios are treated equivalently. The extent of missing data also varies depending on the sensor (Figure 2) and the country where the experiment was conducted (e.g., 60% of Mongolian sensor data is missing). To address the issue of missing data, the first step involved discarding sensor groups with over 90% missing data, aligning with prior work (Santani et al., 2018). This exclusion specifically applied to the ‘cellular_gsm’ features (refer to Figure 2). Subsequently, the remaining missing observations were estimated using the k-nearest-neighbor (kNN) algorithm (Zhou et al., 2018; Xu et al., 2021), with an experimentally determined optimal value of two for k. Following the procedure outlined in prior research (Assi et al., 2023; Zhou et al., 2018; Xu et al., 2021), estimation was performed on the training set, and the learned model was then employed to impute missing data in the testing sets. To retain information that might be encoded in the missing entries of a given sensor group, binary markers were used to denote that the data had been interpolated (Little and Rubin, 2019).
@@ 5.1. RQ1: Situational and Contextual Aspects Around Social Contexts
This figure depicts the distribution of social contexts across countries as a ratio of total self-reports. Mongolia stands out with more ’with relatives’ self-reports, in contrast to other countries where ’alone’ is the dominant context.
This figure presents the distribution of social context according to the hour of the day for all participants. It shows that self-reports of being alone peak around 8 am and dip at around 8 pm, while reports of being with relatives peak at around 8 pm.
@@
The figure compares ’alone’ and ’not alone’ self-reports based on participants’ valence, semantic location, and concurrent activity. It reveals a higher proportion of ’not alone’ reports associated with positive valence and more ’alone’ reports in other valences. ’Alone’ reports are more frequent at home, while ’not alone’ reports dominate in other locations. ’not alone’ reports have the highest percentage during social activities, and ’alone’ reports have the highest percentage during studying.
@@
Figure 5c depicts the co-occurrence of specific moods (positive and negative valence), activities, and semantic locations with the social context (alone or not). Regarding concurrent activities and whether someone is alone or not, activities such as studying, listening to lectures, sleeping, and personal care were predominantly carried out alone. In contrast, activities involving social interaction, such as family care, eating, and nurturing one’s social life, were typically done with others. Furthermore, regarding semantic location, participants reported having been alone at home on 125,270 instances, which could be influenced by the COVID-19 pandemic and the prevalence of remote work and study environments (Fried et al., 2022). However, given the remote work culture that is prevalent even after the pandemic, we could expect these patterns to hold to an extent (Lund et al., 2021; Broom, 2021; Mass, 2021). In terms of gender-specific differences, there were no notable disparities between men and women. Men reported being alone in 54.6% of all self-reports, while women reported to be slightly less alone, accounting for 49.1% of all self-reports. Additionally, women reported to have been with their families, averaging 31.6%, which is approximately 5% more than men. All other distinctions between women and men were minor in the broader context. These findings align with existing research indicating that young men tend to experience more loneliness (Barreto et al., 2021) and tend to be alone more in certain contexts (Fry and Parker, 2021). Prior work also discussed similar findings that women tend to seek more social ties compared to men (Tamminen et al., 2019). However, it is important to note that being alone could be associated with a wide range of factors, including culture, geography, age, etc., in addition to gender (Barreto et al., 2021).
@@
[…]
This table lists the top ten sensing features for discerning ’alone’ or ’not alone’ social contexts, based on p-values after Bonferroni correction from mixed effects models, for individual countries and the dataset as a whole. It shows distinct patterns of app usage and activity types in different countries, such as negative associations of app usage with being alone in Denmark, contrasting with the patterns in Paraguay and Italy. feature coefficient std. error z-score p-value feature coefficient std. error z-score p-value Denmark app_news_and_magazines -0.105 0.041 -2.537 0.011 Paraguay app_finance -0.043 0.012 -3.468 0.000 app_casual -0.041 0.016 -2.492 0.012 notifications_w/o_duplicates -0.018 0.005 -3.248 0.001 app_simulation -0.180 0.074 -2.416 0.015 app_productivity -0.000 $0$3.195 0.001 app_action -0.078 0.032 -0.241 0.015 location_altitude -0.000 $0$3.175 0.001 app_books_and_reference -2.651 1.099 -2.412 0.015 activity_in_vehicle -0.004 0.001 -2.807 0.004 app_sports -0.582 0.241 -2.412 0.015 activity_running -0.025 0.010 -2.588 0.009 app_adventure -0.109 0.045 -2.409 0.015 notifications_removed -0.003 0.001 -2.577 0.009 proximity_std -0.027 0.011 -2.375 0.017 wifi_number_of_devices -0.007 0.003 -2.466 0.013 app_business -0.017 0.007 -2.364 0.018 activity_on_bicycle -0.016 0.006 -2.405 0.016 app_maps_and_navigation -0.352 0.149 -2.354 0.018 wifi_minimum_rssi -0.006 0.002 -2.172 0.029 UK user_presense_time -0.000 0.000 1.987 0.046 Mongolia app_video_players_editors 0.000 0.000 2.613 0.008 proximity_minimum -0.008 0.004 -1.975 0.048 app_shopping -0.056 0.021 -2.609 0.009 proximity_mean -0.018 0.009 -1.909 0.056 app_travel_and_local -0.056 0.022 -2.500 0.012 screen_number_of_episodes 0.011 0.006 -1.844 0.065 app_trivia -0.674 0.284 -2.373 0.017 app_photography 0.006 0.003 1.802 0.701 app_books_and_reference -0.016 0.007 -2.312 0.020 steps_counter 0.000 0.000 1.775 0.075 app_educational -0.109 0.047 -2.300 0.021 proximity_maximum -0.015 0.009 -1.608 0.107 app_weather -3.481 1.516 -2.295 0.021 app_entertainment 0.020 0.012 1.588 0.112 app_comics -0.379 0.165 -2.291 0.021 activity_on_bicycle 0.007 0.004 1.559 0.118 app_art_and_design -0.705 0.315 -2.238 0.025 app_business 0.003 0.002 1.440 0.149 app_news_and_magazines -0.134 0.060 -2.227 0.025 Italy activity_in_vehicle 0.004 0.000 5.970 0.000 All app_news_and_magazines -0.105 0.041 -2.537 0.011 activity_onfoot 0.003 0.000 4.092 0.000 app_casual -0.041 0.016 -2.492 0.012 activity_walking 0.003 0.000 4.010 0.000 app_simulation -0.180 0.074 -2.416 0.015 steps_counter 0.000 0.000 3.243 0.001 app_action -0.078 0.032 -2.414 0.015 activity_tilting 0.000 0.000 3.225 0.001 app_books_and_reference -2.651 1.099 -2.412 0.015 location_altitude 0.000 0.000 3.164 0.001 app_sports -0.582 0.241 -2.412 0.015 proximity_mean -0.013 0.004 -3.067 0.002 app_adeventure -0.109 0.045 -2.409 0.015 notifications_w/o_duplicates -0.006 0.002 -2.409 0.015 proximity_std -0.027 0.011 -2.375 0.017 wifi_std_rssi 0.004 0.001 2.394 0.016 app_business -0.017 0.007 -2.364 0.018 screen_number_of_episodes 0.003 0.001 2.111 0.034 app_maps_and_navigation -0.352 0.149 -2.354 0.018
@@
The statistical analysis presented in Table 3 highlights the top ten sensing features effective in discerning whether a person is alone or not, based on the p-value after Bonferroni correction. This analysis accounts for both fixed (features) and random effects (study participants), offering a nuanced understanding of the data. In Denmark, app usage features such as news and magazines, simulation, and books and references show moderate to high negative coefficients, indicating that using these apps is inversely related to being with others. Conversely, in Paraguay, app usage features like finance and productivity, along with the number of notifications received, have smaller negative coefficients. In the UK, features related to user presence and proximity are statistically significant but show almost no relationship with social context. Similarly, in Mongolia, features like trivia, comics, weather, and arts and design have notable negative relationships with being alone. Italy displays a unique pattern where physical activities such as being in a vehicle, on foot, and walking, though statistically significant, have very small coefficients, showing no clear association. This pattern, at least in terms of feature set, is akin to the UK and Paraguay but contrasts with Denmark and Mongolia, where app usage features have the top statistical significance values. It is also worth noting that, in the UK, except for two features (user_presense_time and proximity_minimum), other features had $p\geq 0.05$, suggesting no statistical significance. The ‘All’ category, aggregating data from all countries, indicates that app features like news and magazines, books and references, maps, sports, and simulation are generally negatively associated with being alone. These findings highlight country-specific differences in the relationship between sensor data features and social context, emphasizing the need to consider user behavior diversity in different countries for understanding and inferring everyday life behaviors with multimodal sensor data (Khwaja et al., 2019; Assi et al., 2023).